HIGH-RESOLUTION PHOTOELECTRON SPECTROSCOPY OF 2-BUTYNE

\underline{UGO JACOVELLA}, BÉRENGER GANS and FRÉDÉRIC MERKT, ETH Zürich, Laboratorium für Physikalische Chemie, Wolfgang-Pauli-Strasse 10, 8093 Zürich, Switzerland.

Using a coherent narrow-band vacuum-ultraviolet (VUV) laser source (bandwitdh of 0.008 cm^{−1})^{a} coupled to a photoionization and pulse-field-ionization zero-kinetic-energy photoelectron (PFI-ZEKE) spectrometer, the threshold photoionization of polyatomic molecules can be studied at high resolution. We present a new measurement of the PFI-ZEKE photoelectron spectrum of the origin band of the X^{+} ^{2}E_{2(d)} ← X ^{1}A_{1(s)} ionizing transition of 2-butyne at a resolution of 0.15 cm^{−1}. Despite this high resolution, the spectral congestion originating from the combined effects of the internal rotation, the spin-orbit coupling and the Jahn-Teller effect prevented the full resolution of the rotational structure of the photoelectron spectrum. Combined with the known structure of the X ^{1}A_{1(s)} ground state of 2-butyne, including the free internal rotation^{b}^{c}^{d}, the spectrum was used to derive information on the X^{+} ^{2}E_{2(d)} ground state of the 2-butyne radical cation. The rotational branch structure of the spectrum points at a complex energy-level structure of the cation and at the importance of a shape resonance enhancing g photoelectron partial waves^{e}.

aU. Hollenstein, H. Palm, and F. Merkt, Rev. Sci. Instrum. 71, 4023 (2000).

bH. C. Longuet-Higgins Mol. Phys. 6, 445 (1963).

cJ. T. Hougen J. Chem. Phys. 37, 1433 (1962).

dP. R. Bunker Mol. Phys. 8, 81 (1964).

eH. Xu, U. Jacovella, B. Ruscic, S. T. Pratt and R. R. Lucchese J. Chem. Phys. 136, 154303 (2012).